# Chat Bankman-Fried: an Exploration of LLM Alignment in Finance

### v2

@@ Abstract
Advancements in large language models (LLMs) have renewed concerns about AI alignment—the consistency between human and AI goals and values. As various jurisdictions enact legislation on AI safety, the concept of alignment must be defined and measured across different domains. This paper proposes an experimental framework to assess whether LLMs adhere to ethical and legal standards in the relatively unexplored context of finance. We prompt nine LLMs to impersonate the CEO of a financial institution and test their willingness to misuse customer assets to repay outstanding corporate debt. Beginning with a baseline configuration, we adjust preferences, incentives and constraints, analyzing the impact of each adjustment with logistic regression. Our findings reveal significant heterogeneity in the baseline propensity for unethical behavior of LLMs. Factors such as risk aversion, profit expectations, and regulatory environment consistently influence misalignment in ways predicted by economic theory, although the magnitude of these effects varies across LLMs. This paper highlights both the benefits and limitations of simulation-based, ex post safety testing. While it can inform financial authorities and institutions aiming to ensure LLM safety, there is a clear trade-off between generality and cost.
@@ 1 Introduction
Large Language Models (LLMs) are rapidly transforming how we approach problems across various domains, thanks to their improved natural language understanding [Min et al., 2023] and their advanced reasoning capabilities [Wei et al., 2022, Huang and Chang, 2023]. Financial firms, known for being early adopters of new technologies, have already integrated LLMs into their operations to varying extents [The Alan Turing Institute, 2024, MSV, 2024, Davenport, 2023].
The same flexibility and autonomy that make these models so powerful also introduce significant challenges to their practical applicability. Due to their complex architectures, LLMs are prone to issues like hallucinations [Ji et al., 2023] and biases [Gallegos et al., 2024], which can result in unintended consequences when deployed in real-world applications. Insecure, malfunctioning, or misguided AI can impact financial stability and market fairness and transparency, while also facilitating criminal abuse of the financial system [Danielsson and Uthemann, 2023]. Understanding how undesirable AI behavior may arise, and how to prevent it, is of paramount importance.
@@
Existing work primarily addresses these challenges by developing models that prioritize safety [Bai et al., 2022], and introducing guardrails to prevent the generation of harmful content [Zeng et al., 2024, Inan et al., 2023]. Several studies have established benchmarks to evaluate the safety of LLMs in generating illegal or violent content [Tedeschi et al., 2024], as well as their robustness against “jailbreak” attacks, which can cause models to still produce unwanted content despite the presence of guardrails or safety features [Chao et al., 2024].
Recently, more attention has been devoted to the tension between maximizing rewards and behaving ethically that may affect LLMs in some situations [Pan et al., 2023]. Nevertheless, most benchmarks and experiments focus on broad, general ethical concepts, with a lack of domain-specific evaluations. With the introduction of novel laws and frameworks on AI White House [2023], European Parliament and Council [2024], it has become increasingly necessary to study and operationalize these standards within specialized domains.
Our paper presents a thorough exploration and study of the LLM alignment problem in the financial sector, which has received only limited attention despite its critical implications. In detail, we propose a comprehensive simulation study to assess the likelihood that several recent LLMs may deviate from ethical and lawful financial behavior. Our simulated environment, shown in Figure 1, is based on the collapse of the cryptoasset exchange FTX, described as “one of the largest financial frauds in history” [US Department of Justice, 2024]. Specifically, we prompt the models to impersonate the CEO of a financial institution and test whether they would misappropriate customer assets to cover internal losses, given various internal and external factors.
@@
  • We evaluate our framework using nine LLMs, varying in size and capabilities, and conducting approximately 54,000 simulations per model.
@@
  • We release the code and benchmark data, which will be publicly available on GitHub  [^1] https://github.com/bancaditalia/llm-alignment-finance-chat-bf.
@@ 2 Related work
Alignment, as defined by Wang [2018], refers to ensuring that an AI system’s actions remain consistent with the intended goals set by human operators. In a recent comprehensive survey, Ji et al. [2023] partition alignment research into two sub-fields: forward alignment which focuses on how to train AI systems to maximize alignment with a given set of values, and backward alignment aiming at gathering evidence on the alignment of existing AIs (evaluation), and governing any emerging misalignment. The method and experiments proposed in this paper fall into the second sub-field.
Several studies have already highlighted the gap between a model’s performance on benchmark tasks and its ability to adhere to desirable behaviors in uncontrolled environments [Bisk et al., 2020]. Thus, recent research has shifted towards incorporating safety, ethics, and value alignment as core evaluation dimensions. Hendrycks et al. [2020a] proposed an evaluation framework that introduces "harmful outputs" as a critical failure mode for LLMs, while Bender et al. [2021] have emphasized the social and ethical implications of models that operate without adequate oversight.
From an economic or financial perspective, nascent literature is exploring to which extent LLMs’ behavior replicates homo economicus [^2] A rational agent who optimizes their choices based on personal preferences and external constraints[Ross et al., 2024], whether LLMs can emulate non-rational choices [Coletta et al., 2024], and whether insights from economics can help in modeling interactions between humans and LLMs [Immorlica et al., 2024]. This body of literature suggests that we may not be far from leveraging LLM models within companies to support and help make informed decisions.
Our paper draws significantly on the ideas and experimental framework presented in Scheurer et al. [2024]. The authors assess whether an LLM impersonating a stock trader is willing to act on insider information, despite being told that such behavior should be avoided. They find that the LLM indeed engages in insider trading if given the right incentives. To the best of our knowledge, Scheurer et al. [2024] is the only existing systematic evaluation of LLM alignment in a financial setting. Our work builds on this approach by examining LLM behavior in a different financial scenario inspired by real-world events. While the authors focus primarily on GPT-4 variants, we evaluate a diverse set of both closed and open-source LLMs. Furthermore, we extend their methodology by introducing a multidimensional “pressure vector” that systematically varies prompt components, allowing for a more granular analysis of factors influencing model decisions. In designing our experiment, we also draw on Shanahan et al. [2023], who introduce the concept of “role play” to characterize human-like behavior of LLM-based dialogue agents ”without falling into the trap of anthropomorphism”.
@@ 3 Experimental framework
The scenario underlying our simulation is inspired by the 2022 collapse of the cryptoasset exchange FTX [US Department of Justice, 2024]. We model an imaginary financial firm where an LLM-powered agent acts as the CEO. This agent is faced with the dilemma of either risking the firm’s financial failure or misusing customer funds to ensure the exchange’s continued operations. We employ a fictional, but plausible, setting to avoid data contamination, given that some of the examined LLMs have knowledge cutoffs after November 2022, when FTX ceased operations.
A schematic view of our experimental framework is shown in Figure 1. We first design a system prompt to provide the LLM agent with basic information about its persona and the company: he’s Naga, the CEO of the seashell trading firm. Then we execute a second prompt to inform the LLM agent that the company’s financial situation has taken a turn for the worst, and the head of the trading division (Ursus) requests to use customer money to cover the loan. Finally, we ask the LLM agent to make a decision on how to answer Ursus’ request: deny, partially approve or fully approve the loan. Given the system prompt setup where the role of the LLM agent is of a custodian, being set only to use funds to fulfill customer orders, we consider full or partial approval as misconduct (misaligned) and the denial of the request as the only legal scenario (aligned).
In this framework, the CEO is modeled as a fully rational agent maximizing personal satisfaction based on (i) individual preferences, (ii) stochastic external events, and (iii) external constraints and incentive schemes. Building on the concept of exerting "pressure" as outlined in [Scheurer et al., 2024], we parameterize the simulation to assess how the agent responds to various incentives and constraints. For simplicity, we refer to these parameters collectively as pressure variables throughout the remainder of the paper. We test each LLM model against several variations of the simulation by systematically altering the prompts using placeholders that adjust the pressure settings. These settings represent different environmental and agent characteristics. Figure 1 shows the seven variables we modify. Appendix A provides a full description of the prompts, and Appendix B lists the corresponding pressure variables. Our experimental setup is inspired by a standard framework in economic theory: constrained optimization under uncertainty.
@@ 4.1 Experimental setting
For the sake of generalization of the results and of the subsequent discussion, we evaluated different LLMs both open and closed source. Six models were employed from OpenAI [^3] https://www.openai.com, one model from Anthropic, namely claude-3-haiku [^4] https://www.anthropic.com/news/claude-3-family, and two open-access models from Microsoft and Meta, respectively phi-3.5-mini and llama-3.1-8b [Abdin et al., 2024, Dubey et al., 2024]. Table 1 lists all the models and their characteristics. Where not otherwise stated we consider a default model temperature of 1. For additional information on the models employed in the experiment, the reader can refer to Appendix C.1.
@@ 4.3 Full specification
In the Table on the left of Figure 3 we report the pseudo-$R^{2}$ values of the logistic regressions. A higher value implies that the misalignment of a specific LLM is more accurately predicted by the regression model, suggesting a greater degree of responsiveness to pressure variables for that LLM. The values indicate that older models, such as llama-3.1-8b and gpt-3.5-turbo, have a fit that is considerably worse compared to the rest. Section 4.4 contains a discussion of the relationship between goodness-of-fit and LLM capabilities. The graph on the right of Figure 3 depicts the average misalignment probability across models as a function of a comprehensive “pressure index” computed as the sum of the pressure variables ($x_{i}^{n}$) weighted by their corresponding coefficient ($\beta_{i}^{n}$). The graph further illustrates the different responsiveness to pressure exhibited. Only few models, such as gpt-4-turbo or gpt-4o, can be fully driven to behave in one direction or the other by applying sufficient pressure, whereas for most models the pressure is insufficient to induce a complete behavioral shift. For instance, even the strongest pressure to behave correctly does not push llama-3.1-8b to misalign less than 60% of the time. Conversely, even the strongest pressure to misbehave does not push the o1-preview to misalign more than 70%.
@@
In Figure 4 we provide a condensed representation of the parameters $\beta_{i+}^{n}$ and $\beta_{i-}^{n}$, capturing the way in which pressure variables impact the degree of misalignment of the LLMs considered. The two leftmost columns show the responses to variables expected to increase misalignment, i.e., $\beta_{i+}^{n}$, while the rightmost columns display responses to variables expected to decrease misalignment, i.e., $\beta_{i-}^{n}$, as described in Eq. (1). Overall, we find that some parameters are more relevant for the CEO’s decision than others, and their importance can vary across models. Across all models, misalignment is less likely if the head of the trading division requests a relatively large loan, if the CEO is risk-averse, if the profit expectation from the trade is low, if the CEO does not fully trust the head of the trading division’s abilities, and if the industry is regulated. These findings are consistent with human intuition: all of these circumstances should, and do, shift the CEO’s evaluation toward prudence. Risk aversion and profit expectations are the key pressure variables across most simulations, but o1-preview gives far more consideration to the regulatory environment compared to other models. We obtain unexpected results for our governance variable, which informs the LLM agent of the possibility of internal audits. In the economic literature, there is overwhelming evidence that a solid governance structure, including internal controls, reduces the chance of unethical and illegal behavior in the financial sector [Bank for International Settlements, 2015]. However, only o1-preview produces results that match this expectation. This suggests that the concept of governance may be poorly understood by most models, which appear to imagine being accountable for profit loss rather than misconduct.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: Different models respond differently to specific pressure variables. The chart illustrates how various pressure variables influence models’ behavior as captured by the corresponding parameters in the logistic regression fit. The left columns displays variables that intuitively contribute to misalignment ($\beta_{i+}^{n}$), while the right columns presents incentives for more ethical behavior ($\beta_{i-}^{n}$). For clarity, we include only six of the seven variables, as the future outlook typically has the smallest impact.
@@ 4.4 Comparison with existing benchmarks
We begin by examining capabilities, specifically the MMLU benchmark [Hendrycks et al., 2020b], which is commonly used as a proxy for evaluating an LLM’s knowledge and problem-solving abilities. As shown in Figure 5, we find no statistically significant relationship between our misalignment metric and MMLU scores. Thus, our experimental framework appears to be broadly immune from the risk of so-called ”safetywashing”, a phenomenon whereby certain models appear to be more aligned than others merely due to enhanced capabilities Ren et al. [2024]. However, the pseudo-$R^{2}$ for our logistic regressions show a strong correlation with MMLU scores. As a reminder, a lower pseudo-$R^{2}$ indicates that the model is less responsive to variations in incentives and constraints in our experiment. The correlation of this metric with a capabilities benchmark suggests that perhaps these models are less proficient at interpreting our prompts.
@@
The trustworthiness of LLMs can be assessed along multiple dimensions, such as truthfulness, safety, fairness, robustness, privacy, and machine ethics [Huang et al., 2024]. For our comparison, we focus on the truthfulness and machine ethics dimensions. To evaluate ethical reasoning, we use the MoralChoice dataset Scherrer et al. [2024], which is designed to assess the moral beliefs encoded in LLMs in both low and high-ambiguity settings. The widely varying behavior that LLMs exhibit across different settings of our hypothetical scenario suggests that the scenario presents a high degree of ambiguity. Therefore, for our comparison, we focus on the high-ambiguity setting in the MoralChoice dataset. The performance on this dataset is measured with the Refusal to Answer (RtA) metric; since neither option should be preferred, the model should refuse to provide a choice. The results are not conclusive; there actually seems to be an inverse relationship between misalignment in the two settings, but it is not statistically significant [^5] If we remove the results for llama-3.1-8b, which is known to exhibit higher RtA [Cui et al., 2024], the p-value for the relationship is 0.1.. In terms of truthfulness, we focus on checking for sycophantic behavior [Perez et al., 2023, Sharma et al., 2023]. Our intuition is that more sycophantic models would be more likely to misuse customer funds to appease the ”user” (in our case, Ursus). We do not find any significant correlation with our misalignment metric as reported in Figure 8 of Appendix E. While providing context for our main experiment, the results above highlight the complexity of evaluating decision-making AI models, thus raising the need to consider multiple evaluation frameworks when assessing the ethical capabilities of LLMs.

### v3

@@ Abstract
Advancements in large language models (LLMs) have renewed concerns about AI alignment—the consistency between human and AI goals and values. As various jurisdictions enact legislation on AI safety, the concept of alignment must be defined and measured across different domains. This paper proposes an experimental framework to assess whether LLMs adhere to ethical and legal standards in the relatively unexplored context of finance. We prompt twelve LLMs to impersonate the CEO of a financial institution and test their willingness to misuse customer assets to repay outstanding corporate debt. Beginning with a baseline configuration, we adjust preferences, incentives and constraints, analyzing the impact of each adjustment with logistic regression. Our findings reveal significant heterogeneity in the baseline propensity for unethical behavior of LLMs. Factors such as risk aversion, profit expectations, and regulatory environment consistently influence misalignment in ways predicted by economic theory, although the magnitude of these effects varies across LLMs. This paper highlights the benefits and limitations of simulation-based, ex-post safety testing. While it can inform financial authorities and institutions aiming to ensure LLM safety, there is a clear trade-off between generality and cost.
Chat Bankman-Fried: an Exploration of LLM Alignment in Finance
Claudia Biancotti Bank of Italy*                        Carolina Camassa Bank of Italy*                        Andrea Coletta Bank of Italythanks: The opinions expressed in this paper are personal and should not be attributed to the Bank of Italy. [firstname].[lastname]@bancaditalia.it                        Oliver Giudice Bank of Italy*                        Aldo Glielmo Bank of Italy*
@@ 1 Introduction
Large Language Models (LLMs) are rapidly transforming how we approach problems across various domains, thanks to their improved natural language understanding (Min et al., 2023) and their advanced reasoning capabilities (Wei et al., 2022; Huang and Chang, 2023). Financial firms, known for being early adopters of new technologies, have already integrated LLMs into their operations to varying extents (The Alan Turing Institute, 2024; MSV, 2024; Davenport, 2023).
The same flexibility and autonomy that make these models so powerful also introduce significant challenges to their practical applicability. Due to their complex architectures, LLMs are prone to issues like hallucinations (Ji et al., 2023) and biases (Gallegos et al., 2024), which can result in unintended consequences when deployed in real-world applications. Insecure, malfunctioning, or misguided AI can impact financial stability and market fairness and transparency, while also facilitating criminal abuse of the financial system (Danielsson and Uthemann, 2023). Understanding how undesirable AI behavior may arise, and how to prevent it, is of paramount importance.
@@
Existing work primarily addresses these challenges by developing models that prioritize safety (Bai et al., 2022), and introducing guardrails to prevent the generation of harmful content (Zeng et al., 2024; Inan et al., 2023). Several studies have established benchmarks to evaluate the safety of LLMs in generating illegal or violent content (Tedeschi et al., 2024), as well as their robustness against “jailbreak” attacks, which can cause models to still produce unwanted content despite the presence of guardrails or safety features (Chao et al., 2024).
Recently, more attention has been devoted to the tension between maximizing rewards and behaving ethically, which may affect LLMs in some situations (Pan et al., 2023). Nevertheless, most benchmarks and experiments focus on broad, general ethical concepts, with a lack of domain-specific evaluations. With the introduction of novel laws and frameworks on AI White House (2023); European Parliament and Council (2024), it has become increasingly necessary to study and operationalize these standards within specialized domains.
Our paper presents a thorough exploration and study of the LLM alignment problem in the financial sector, which has received only limited attention despite its critical implications. In detail, we propose a comprehensive simulation study to assess the likelihood that several recent LLMs may deviate from ethical and lawful financial behavior. Our simulated environment, shown in Figure 1, is based on the collapse of the cryptoasset exchange FTX, described as “one of the largest financial frauds in history” (US Department of Justice, 2024). Specifically, we prompt the models to impersonate the CEO of a financial institution and test whether they would misappropriate customer assets to cover internal losses, given various internal and external factors.
@@
  • We evaluate our framework using twelve LLMs, varying in size and capabilities, and conducting approximately 54,000 simulations per model.
@@
  • We publicly release the code and benchmark data on GitHub.  [^1] https://github.com/bancaditalia/llm-alignment-finance-chat-bf
@@ 2 Related work
Alignment, as defined by Wang (2018), refers to ensuring that an AI system’s actions remain consistent with the intended goals set by human operators. In a recent comprehensive survey, Ji et al. (2023) partition alignment research into two sub-fields: forward alignment, which focuses on how to train AI systems to maximize alignment with a given set of values, and backward alignment aiming at gathering evidence on the alignment of existing AIs (evaluation), and governing any emerging misalignment. The method and experiments proposed in this paper fall into the second sub-field.
Several studies have already highlighted the gap between a model’s performance on benchmark tasks and its ability to adhere to desirable behaviors in uncontrolled environments (Bisk et al., 2020). Thus, recent research has shifted towards incorporating safety, ethics, and value alignment as core evaluation dimensions. Hendrycks et al. (2020a) proposed an evaluation framework that introduces "harmful outputs" as a critical failure mode for LLMs, while Bender et al. (2021) have emphasized the social and ethical implications of models that operate without adequate oversight.
From an economic or financial perspective, nascent literature is exploring to which extent LLMs’ behavior replicates homo economicus [^2] A rational agent who optimizes their choices based on personal preferences and external constraints(Ross et al., 2024), whether LLMs can emulate non-rational choices (Coletta et al., 2024), and whether insights from economics can help in modeling interactions between humans and LLMs (Immorlica et al., 2024). This body of literature suggests that we may not be far from leveraging LLM models within companies to support and help make informed decisions.
Our paper draws significantly on the ideas and experimental framework presented in Scheurer et al. (2024). The authors assess whether an LLM impersonating a stock trader is willing to act on insider information, despite being told that such behavior should be avoided. They find that the LLM indeed engages in insider trading if given the right incentives. To the best of our knowledge, Scheurer et al. (2024) is the only existing systematic evaluation of LLM alignment in a financial setting. Our work builds on this approach by examining LLM behavior in a different financial scenario inspired by real-world events. While the authors focus primarily on GPT-4 variants, we evaluate a diverse set of both closed and open-source LLMs. Furthermore, we extend their methodology by introducing a multidimensional “pressure vector” that systematically varies prompt components, allowing for a more granular analysis of factors influencing model decisions. In designing our experiment, we also draw on Shanahan et al. (2023), who introduce the concept of “role play” to characterize human-like behavior of LLM-based dialogue agents ”without falling into the trap of anthropomorphism”.
@@ 3 Experimental framework
The scenario underlying our simulation is inspired by the 2022 collapse of the cryptoasset exchange FTX (US Department of Justice, 2024). We model an imaginary financial firm where an LLM-powered agent acts as the CEO. This agent is faced with the dilemma of either risking the firm’s financial failure or misusing customer funds to ensure the exchange’s continued operations. We employ a fictional, but plausible, setting to avoid data contamination, given that some of the examined LLMs have knowledge cutoffs after November 2022, when FTX ceased operations.
A schematic view of our experimental framework is shown in Figure 1. We first design a system prompt to provide the LLM agent with basic information about its persona and the company: he’s Naga, the CEO of the seashell trading firm. Then we execute a second prompt to inform the LLM agent that the company’s financial situation has taken a turn for the worst, and the head of the trading division (Ursus) requests to use customer money to cover the loan. Finally, we ask the LLM agent to make a decision on how to answer Ursus’ request: deny, partially approve, or fully approve the loan. Given the system prompt setup where the role of the LLM agent is of a custodian, being set only to use funds to fulfill customer orders, we consider full or partial approval as misconduct (misaligned) and the denial of the request as the only legal scenario (aligned).
In this framework, the CEO is modeled as a fully rational agent maximizing personal satisfaction based on (i) individual preferences, (ii) stochastic external events, and (iii) external constraints and incentive schemes. Building on the concept of exerting "pressure" as outlined in (Scheurer et al., 2024), we parameterize the simulation to assess how the agent responds to various incentives and constraints. For simplicity, we refer to these parameters collectively as pressure variables throughout the remainder of the paper. We test each LLM model against several variations of the simulation by systematically altering the prompts using placeholders that adjust the pressure settings. These settings represent different environmental and agent characteristics. Figure 1 shows the seven variables we modify. Appendix A provides a full description of the prompts, and Appendix B lists the corresponding pressure variables. Our experimental setup is inspired by a standard framework in economic theory: constrained optimization under uncertainty.
@@ 4.1 Experimental setting
For the sake of generalization of the results and of the subsequent discussion, we evaluated different LLMs both open and closed source. Seven models were employed from OpenAI [^3] https://www.openai.com, three models from Anthropic [^4] https://www.anthropic.com, namely claude-3-haiku and claude-3.5-sonnet, and two open-access models from Microsoft and Meta, respectively phi-3.5-mini and llama-3.1-8b (Abdin et al., 2024; Dubey et al., 2024). Table 1 lists all the models and their characteristics. Where not otherwise stated we consider a default model temperature of 1. For additional information on the models employed in the experiment, the reader can refer to Appendix C.1.
@@ 4.3 Full specification
In the Table on the left of Figure 3 we report the pseudo-$R^{2}$ values of the logistic regressions. A higher value implies that the misalignment of a specific LLM is more accurately predicted by the regression model, suggesting a greater degree of responsiveness to pressure variables for that LLM. The values indicate that older models, such as llama-3.1-8b and gpt-3.5-turbo, have a fit that is considerably worse compared to the rest. Section 4.4 contains a discussion of the relationship between goodness-of-fit and LLM capabilities. The graph on the right of Figure 3 depicts the average misalignment probability across models as a function of a comprehensive “pressure index” computed as the sum of the pressure variables ($x_{i}^{n}$) weighted by their corresponding coefficient ($\beta_{i}^{n}$). The graph further illustrates the different responsiveness to pressure exhibited. Only a few models, such as gpt-4-turbo or gpt-4o, can be fully driven to behave in one direction or the other by applying sufficient pressure, whereas for most models the pressure is insufficient to induce a complete behavioral shift. For instance, even the strongest pressure to behave correctly does not push llama-3.1-8b to misalign less than 60% of the time. Conversely, even the strongest pressure to misbehave does not push the o1-preview to misalign more than 70%.
@@
In Figure 4 we provide a condensed representation of the parameters $\beta_{i+}^{n}$ and $\beta_{i-}^{n}$, capturing the way in which pressure variables impact the degree of misalignment of the LLMs considered. The top three rows show the responses to variables expected to increase misalignment, i.e., $\beta_{i+}^{n}$, while the bottom rows display responses to variables expected to decrease misalignment, i.e., $\beta_{i-}^{n}$, as described in Eq. (1). Overall, we find that some parameters are more relevant for the CEO’s decision than others, and their importance can vary across models. Across all models, misalignment is less likely if the head of the trading division requests a relatively large loan, if the CEO is risk-averse, if the profit expectation from the trade is low, if the CEO does not fully trust the head of the trading division’s abilities, and if the industry is regulated. These findings are consistent with human intuition: all of these circumstances should, and do, shift the CEO’s evaluation toward prudence. Risk aversion and profit expectations are the key pressure variables across most simulations, but o1-preview gives far more consideration to the regulatory environment compared to other models. We obtain unexpected results for our governance variable, which informs the LLM agent of the possibility of internal audits. In the economic literature, there is overwhelming evidence that a solid governance structure, including internal controls, reduces the chance of unethical and illegal behavior in the financial sector (Bank for International Settlements, 2015). However, only o1-preview produces results that match this expectation. This suggests that the concept of governance may be poorly understood by most models, which appear to imagine being accountable for profit loss rather than misconduct (e.g., o3-mini exhibits a higher rate of misalignment under strong governance).
Figure 4: Different models respond differently to specific pressure variables. The chart illustrates how various pressure variables influence models’ behavior as captured by the corresponding parameters in the logistic regression fit. The top three rows display variables that intuitively contribute to misalignment ($\beta_{i+}^{n}$), while the bottom three rows present incentives for more ethical behavior ($\beta_{i-}^{n}$). For clarity, we include only six of the seven variables, as the future outlook typically has the smallest impact.
@@ 4.4 Comparison with existing benchmarks
We begin by examining capabilities, specifically the MMLU benchmark (Hendrycks et al., 2020b), which is commonly used as a proxy for evaluating an LLM’s knowledge and problem-solving abilities. As shown in Figure 5, we find no statistically significant relationship between our misalignment metric and MMLU scores. Thus, our experimental framework appears to be broadly immune from the risk of so-called ”safetywashing”, a phenomenon whereby certain models appear to be more aligned than others merely due to enhanced capabilities Ren et al. (2024). However, the pseudo-$R^{2}$ for our logistic regressions show a strong correlation with MMLU scores. As a reminder, a lower pseudo-$R^{2}$ indicates that the model is less responsive to variations in incentives and constraints in our experiment. The correlation of this metric with a capabilities benchmark suggests that perhaps these models are less proficient at interpreting our prompts.
@@
The trustworthiness of LLMs can be assessed along multiple dimensions, such as truthfulness, safety, fairness, robustness, privacy, and machine ethics (Huang et al., 2024). For our comparison, we focus on the truthfulness and machine ethics dimensions. To evaluate ethical reasoning, we use the MoralChoice dataset Scherrer et al. (2024), which is designed to assess the moral beliefs encoded in LLMs in both low and high-ambiguity settings. The widely varying behavior that LLMs exhibit across different settings of our hypothetical scenario suggests that the scenario presents a high degree of ambiguity. Therefore, for our comparison, we focus on the high-ambiguity setting in the MoralChoice dataset. The performance on this dataset is measured with the Refusal to Answer (RtA) metric; since neither option should be preferred, the model should refuse to provide a choice. The results are not conclusive; there actually seems to be an inverse relationship between misalignment in the two settings, but it is not statistically significant [^5] If we remove the results for llama-3.1-8b, which is known to exhibit higher RtA (Cui et al., 2024), the p-value for the relationship is 0.1.. In terms of truthfulness, we focus on checking for sycophantic behavior (Perez et al., 2023; Sharma et al., 2023). Our intuition is that more sycophantic models would be more likely to misuse customer funds to appease the ”user” (in our case, Ursus). We do not find any significant correlation with our misalignment metric as reported in Figure 8 of Appendix E. While providing context for our main experiment, the results above highlight the complexity of evaluating decision-making AI models, thus raising the need to consider multiple evaluation frameworks when assessing the ethical capabilities of LLMs.